OpenStudy (anonymous):

Find the midpoint of the line segment joining the points P1= (3,4) P2=(-2,6)

OpenStudy (anonymous):

Apply the midpoint formula: \[(\frac{ 3+ (-2) }{ 2 }, \frac{ 4 + 6 }{ 2 }) = (\frac{ 1 }{ 2 }, \frac{ 10 }{ 2 })\] So the result is: \[midpoint = (\frac{ 1 }{ 2 }, 5)\]

OpenStudy (anonymous):

Thank you!!!!